# Transfer Matrix Approach for Topological Edge States

### v2

@@ Abstract
We suggest and develop a novel approach for describing topological properties of a periodic system purely from the transfer matrix associated to a unit cell. This approach is employed to obtain a simple pictorial proof of the Zak phase bulk-edge correspondence for any one dimensional system. We apply the transfer matrix method to study some generalisations of the Su-Schrieffer-Heeger (SSH) model, such as the tetramer SSH4 model and a dimerised one-dimensional photonic crystal. We thus obtain simple conditions for the existence of topologically protected edge states in these models.
@@ I Introduction
Topological insulators are commonly associated with robust states, localised at the edges of the sample, and which existence are entirely predicted by the topology of the bulk [1]. As expected, due to their promised robustness, topological edge states currently form an exciting playground for technological applications. For example, topological insulators have been considered as candidate materials to realise qubits for quantum computing [2, 3], as well as efficient lasing and harmonic generation in photonic systems [4, 5].
@@
In this article, we introduce a new method to analyse one-dimensional topological insulators using the transfer matrix. This transfer matrix method offers an approach complementary to the standard Bloch Hamiltonian method. As a preview, a system with $N\times N$ sites per unit cell and $k$-nearest neighbour hoppings would admit a Bloch Hamiltonian of size $N\times N$, while the transfer matrix would have a size of $2k\times 2k$. In addition, while the Bloch Hamiltonian works best with periodic boundary conditions, the transfer matrix method can be naturally equipped with open or closed boundary conditions. A consequence of this is that the edge states appear naturally within the transfer matrix method, so that we may characterise the edge states more explicitly. We thus apply the transfer matrix method to inversion-symmetric models and obtain simple conditions for the appearance of their topologically protected edge states.
@@ II Transfer Matrix of the SSH model
$\begin{pmatrix}\mu_{1}&\gamma_{1}&&&&\\ \gamma_{1}&\mu_{2}&\gamma_{2}&&&\\ &\gamma_{2}&\mu_{1}&\gamma_{1}&&\\ &&\gamma_{1}&\ddots&\ddots&\\ &&&\ddots&\ddots&\gamma_{1}\\ &&&&\gamma_{1}&\mu_{2}\end{pmatrix}\begin{pmatrix}a_{1}\\ a_{2}\\ a_{3}\\ \vdots\\ \vdots\\ a_{2N}\end{pmatrix}=0$ (2)
@@
The unit cell for the matrix is provided in the left of Eq. (3) below. Solving the matrix equation within each unit cell, we obtain the reduced row echelon form which can be written in the block form $(-T,\mathbbm{1})$ as shown in the right side of Eq. (3):
$\begin{pmatrix}\gamma_{2}&\mu_{1}&\gamma_{1}&0\\ 0&\gamma_{1}&\mu_{2}&\gamma_{2}\end{pmatrix}\to\begin{pmatrix}\frac{\gamma_{2}}{\gamma_{1}}&\frac{\mu_{1}}{\gamma_{1}}&1&0\\ -\frac{\mu_{2}}{\gamma_{1}}&-\frac{\mu_{1}\mu_{2}}{\gamma_{1}\gamma_{2}}+\frac{\gamma_{1}}{\gamma_{2}}&0&1\end{pmatrix}.$ (3)
The condition that a solution is in the null-space of this reduced matrix becomes exactly the definition of the transfer matrix:
$\bm{a}_{i+1}=T\bm{a}_{i},$ (4)
where $\bm{a}_{i}=(a_{2i},a_{2i+1})^{T}$ and $a_{2i}$ is the rightmost amplitude of the previous unit cell and $a_{2i+1}$ is the first (leftmost) amplitude of the current unit cell. In this way, for a system of $N$ sites, we have introduced 2 more amplitudes in our transfer matrix description, $a_{0}$ and $a_{N+1}$. These will be important in the next paragraph when we consider boundary conditions. The gaussian elimination process also shows that in general, any one-dimensional periodic dynamical system admits a set of transfer matrices that is unique up to basis changes or symmetry transformations of the model [^1] This freedom is important as any hermitian tridiagonal Hamiltonian admits a gauge transformation to a real-symmetric tridiagonal Hamiltonian. This choice of basis will become important for the arguments in the later sections..
@@
When the chain is semi-infinite, this condition, along with the condition that $T^{k}(^{0}_{1})$ stays bounded in the chain, recovers the exact eigenmodes of the chain. It should be noted that the Bloch theorem no longer applies in the semi-infinite case as there is no translation symmetry in the system. To recover the exact eigenspectrum one needs to consider each mode as a superposition of counter-propagating waves in the infinite chain which fulfills the condition (5).
@@ III Topological Protection in the Transfer Matrix Formalism
Figure 1: (a) Illustration of the SSH model. The open circles denote sites at which the electron can sit, and the lines denote the coupling strength between each site. In the SSH model, the couplings alternate between two values $\gamma_{1}$ and $\gamma_{2}$. (b) The band structure of the SSH model $(\mu_{1}=\mu_{2})$ embedded in $\mathrm{SL}(2,\mathbb{R})$ (3D Iwasawa parameters space). The parameter $k$ is only defined modulo $2\pi$. Each curve represents a Hamiltonian, or equivalently, a family of $T$-matrices $T(E)$ for a different value of $\gamma_{1}/\gamma_{2}$. When the curve touches one of the orange surfaces, an edge state is supported. The decay length of this edge state is given by $|\ln(a)|^{-1}$. The green regions are the allowed bands: Each point on the curve inside the green region corresponds to a bulk propagating mode.
@@ III.1 The $\mathbb{R}$) Picture for Topological Phases
$T=\begin{pmatrix}\cos(k)&\sin(k)\\ -\sin(k)&\cos(k)\end{pmatrix}\begin{pmatrix}\frac{1}{a}&\\ &a\end{pmatrix}\begin{pmatrix}1&\\ n&1\end{pmatrix}$ (6)
where $k,n\in\mathbb{R},a>0$ are the Iwasawa parameters, and they are smooth functions of the entries $T_{ij}$. By doing so, we can parameterise $T(E)$ as a curve in a three-dimensional space, where the coordinates are taken to be the Iwasawa parameters: $k(E)$, $a(E)$, $n(E)$. We plot a family of curves with various choices of $\gamma$ in Fig. 1 for the SSH model. Since $\det T\in\mathrm{SL}(2,\mathbb{R})$ the condition that the bulk states must stay bounded in the chain is automatically satisfied and we may identify the transfer matrices supporting bulk states by simply looking at the condition $\mathrm{Tr}(T)\leq 2$. On the other hand for a $T$ matrix to support an edge state, we require $(0,1)$ to be an eigenvector of $T$ with eigenvalue less than one. So, we need $k\in\pi\mathbb{Z}$ and $a<1$.
@@
Topological insulators are characterised by edge states which cannot disappear unless the material passes through a conducting state [21]. In the above picture, this means that we need to somehow exclude all the points $k=\pi\mathbb{Z},\ln(a)=0,n\neq 0$. Any kind of constraint on the Hamiltonian or $T$-matrix which excludes these points will do the job just fine, but this is usually done using symmetries. In agreement with Zak’s argument [22], parity symmetry automatically gives us
$\sigma_{x}T\sigma_{x}=T^{-1}\quad\implies\quad n=\left(\frac{1}{a^{2}}-1\right)\tan(k)$ (7)
@@ III.2 Proof of the Zak Phase Bulk-Edge Correspondence
Here, we show how our formalism so far is connected to Bloch band topological invariants. In doing so, we will obtain a rigorous but pictorially simple proof of the bulk-edge correspondence. We stress that our results hold for any number of sites per unit cell, as well as for continuous quantum mechanical or photonic systems. This will become more clear in section IV.1 where we treat the case of $4$ sites per unit cell more explicitly. But for now, we will simply highlight a few key features when we increase the number of sites per unit cell.
@@
Here $\mathcal{N}$ is a normalisation constant, $\phi(K)$ is a real function, and the entries in the dots are not important (see Appendix A for details).
@@
$\phi_{\text{Zak}}\equiv\frac{1}{\mathrm{i}\mkern 1.0mu}\int\limits_{0}^{2\pi}\left\langle K\right|\partial_{K}\left|K\right\rangle\mathop{}\!\mathrm{d}K=\frac{\phi(2\pi)-\phi(0)}{2}$ (9)
where $K$ is the Bloch eigenvector. This result is well-known for the case of 2 sites per unit cell [25], but to the best of our knowledge has not been derived in the general case of $N$ sites per unit cell. We provide details of the general case in Appendix A.
@@
$\phi(K)=\mathrm{arg}\left[\frac{\langle 0|K\rangle_{T}}{\langle 1|K\rangle_{T}}\left\langle K\right|_{T}T\left|K\right\rangle_{T}\right]=K-\varphi(K),$ (10)
where $\left|0\right\rangle$ and $\left|1\right\rangle$ are the standard position basis vectors [^4] For a finite chain of $N$ sites with Dirichlet boundary conditions this is defined to be a basis such that the boundary condition is given by $a_{0}=0$ and $a_{N+1}=0$ as in equation (5)., $\left|K\right\rangle_{T}$ is the eigenvector of $T$ with eigenvalue $\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0muK}$, and $\varphi(K)$ is the relative phase of the two entries in the $T$-matrix eigenvector in this basis. It should be noted here that both $\phi$ and $\varphi$ depend on the choice of basis. This is why the middle expression in (10) must be written in a basis-dependent way.
Figure 2: Bulk-edge correspondence for the Zak phase, as viewed in $\mathbb{R}$), and projected to $n$ given by Eq. (7). The paths $T(E)$ for 4-band SSH4 model is plotted for various $\gamma_{13}/\sqrt{\gamma_{2}\gamma_{4}}$ ratios. Note that the parameter $k$ is only defined modulo $2\pi$. The plot has been unfolded in much the same way as a band diagram can be unfolded beyond the first Brillouin zone. The green regions denote transfer matrices which support bulk modes (allowed bands regions), while the brown lines denote transfer matrices which support edge modes. The purple dashed boundaries denote the points where $\phi=\pi$, and the red dashed curves denote the points where $\phi=0$. The Zak phase $\phi_{\text{Zak}}$ is $\pi$ when the $T(E)$ curve enters and exits from boundaries of different colours, and it is 0 when the two colours are the same. As such, we see that the creation of edge states (magenta stars) is necessarily accompanied by the change in Zak phase of the two neighbouring bands.
The parity symmetry condition (7) along with the condition $|\mathrm{Tr}(T)|=2$, that the Bloch eigenvector is $K\equiv 0\pmod{\pi}$, fixes $T$ to lie in one of four branches of solutions. As an example, one of the possible solutions is given by
$T=\left(\begin{array}[]{cc}-\frac{\cos(k)}{\sin(k)+\cos(k)}&\frac{\sin(k)}{\sin(k)+\cos(k)}\\ -\frac{\sin(k)}{\sin(k)+\cos(k)}&-\frac{2\sin(k)+\cos(k)}{\sin(k)+\cos(k)}\\ \end{array}\right),$
@@
This shows that in general the number of edge state in the forbidden zone above the $k^{\text{th}}$ band is the sum of Zak phases of all the lower bands, modulo 2:

### v3

@@ Abstract
We suggest and develop a novel approach for describing topological properties of a periodic system purely from the transfer matrix associated to a unit cell. Our approach uses the Iwasawa decomposition to parametrise the transfer matrix uniquely in terms of three real numbers. This allows us to obtain simple conditions for the existence of topologically protected edge states and to provide a visual illustration of all possible solutions. In order to demonstrate our method in action, we apply it to study some generalisations of the Su-Schrieffer-Heeger (SSH) model, such as the tetramer SSH4 model and a dimerised one-dimensional photonic crystal. Finally, we also obtained a simple pictorial proof of the Zak phase bulk-edge correspondence for any one dimensional system using this approach.
@@ I Introduction
Topological insulators are commonly associated with robust states, localised at the edges of the sample, and which existence are entirely predicted by the topology of the bulk [1]. As expected, due to their promised robustness, topological edge states currently form an exciting playground for technological applications. For example, topological insulators have been considered as candidate materials to realise qubits for quantum computing [2, 3], as well as efficient lasing and harmonic generation in photonic systems [4, 5, 6].
@@
In this article, we introduce a new method to analyse one-dimensional topological insulators based on top of the well-established transfer matrix approach [17, 18, 19]. This transfer matrix method offers an approach complementary to the standard Bloch Hamiltonian method [20, 21]. As a preview, a system with $N$ sites per unit cell and $k$-nearest neighbour hoppings would admit a Bloch Hamiltonian of size $N\times N$, while the transfer matrix would have a size of $2k\times 2k$. In addition, while the Bloch Hamiltonian works best with periodic boundary conditions, the transfer matrix method can be naturally equipped with open or closed boundary conditions. A consequence of this is that the edge states appear naturally within the transfer matrix method, so that we may characterise the edge states more explicitly. We thus apply the transfer matrix method to inversion-symmetric models and obtain simple conditions for the appearance of their topologically protected edge states. This is done by uniquely parametrising the transfer matrix in terms of three real numbers using the Iwasawa decomposition of $\mathrm{SL}(2;\mathbb{R})$. This parametrisation reveals the submanifolds corresponding to symmetry-protected systems, and the topology of these submanifolds is intimately connected with the edge states of the system.
@@ II Transfer Matrix of the SSH model
$\begin{pmatrix}\mu_{1}&\gamma_{1}&&&&\\ \gamma_{1}&\mu_{2}&\gamma_{2}&&&\\ &\gamma_{2}&\mu_{1}&\gamma_{1}&&\\ &&\gamma_{1}&\ddots&\ddots&\\ &&&\ddots&\ddots&\gamma_{1}\\ &&&&\gamma_{1}&\mu_{2}\end{pmatrix}\begin{pmatrix}a_{1}\\ a_{2}\\ a_{3}\\ \vdots\\ \vdots\\ a_{2N}\end{pmatrix}=0,$ (2)
@@
Solving the matrix equation (2) within first unit cell
$\begin{pmatrix}\gamma_{2}&\mu_{1}&\gamma_{1}&0\\ 0&\gamma_{1}&\mu_{2}&\gamma_{2}\end{pmatrix}\begin{pmatrix}a_{0}\\ a_{1}\\ a_{2}\\ a_{3}\end{pmatrix}=0$ (3)
using Gauss-Jordan elimination, we obtain the reduced row echelon form which can be written in the block form:
$\begin{pmatrix}-T&\mathbbm{1}\end{pmatrix}\begin{pmatrix}a_{0}\\ a_{1}\\ a_{2}\\ a_{3}\end{pmatrix}=0,\quad T=\begin{pmatrix}-\frac{\gamma_{2}}{\gamma_{1}}&-\frac{\mu_{1}}{\gamma_{1}}\\ \frac{\mu_{2}}{\gamma_{1}}&\frac{\mu_{1}\mu_{2}}{\gamma_{1}\gamma_{2}}-\frac{\gamma_{1}}{\gamma_{2}}\end{pmatrix},$ (4)
where $\mathbbm{1}$ is a $2\times 2$ identity matrix. The condition that a solution is in the null-space of this reduced matrix becomes exactly the definition of the transfer matrix:
$\bm{a}_{i+1}=T(E)\bm{a}_{i}.$ (5)
Here, the dependence of $T$ on $E$ comes in through the energy-dependence of $\mu_{1},\mu_{2}$ in Eq. (4). Also, we define $\bm{a}_{i}:=(a_{2i},a_{2i+1})^{T}$ where $a_{2i}$ is the rightmost amplitude of the previous unit cell and $a_{2i+1}$ is the first (leftmost) amplitude of the current unit cell. In this way, for a system of $N$ sites, we have introduced 2 more amplitudes in our transfer matrix description, $a_{0}$ and $a_{N+1}$. These will be important in the next paragraph when we consider boundary conditions. The gaussian elimination process also shows that in general, any one-dimensional periodic Hamiltonian admits a set of transfer matrices that is unique up to basis changes or symmetry transformations [^1] This freedom is important as any hermitian tridiagonal Hamiltonian admits a gauge transformation to a real-symmetric tridiagonal Hamiltonian. This choice of basis will become important for the arguments in the later sections..
@@
When the chain is semi-infinite, this condition, along with the condition that $T^{k}(^{0}_{1})$ stays bounded in the chain, recovers the exact eigenmodes of the chain. This includes both the Bloch bands and the edge modes. It should be noted that the Bloch theorem no longer applies in the semi-infinite case as there is no translation symmetry in the system. To recover the exact eigenspectrum one needs to consider each mode as a superposition of counter-propagating waves in the infinite chain which fulfills the condition (6).
In general, for a $T$-matrix with $\det T=1$, the above reasoning reduce to simple conditions. For the bulk modes, for which Bloch’s theorem is applicable, each propagating mode must be an eigenvector of $T$ with eigenvalues $\mathrm{e}\mkern 1.0mu^{\pm\mathrm{i}\mkern 1.0muK}$, where $K$ is the system quasi-momentum. Since the trace of the matrix is equal to the sum of its eigen values, this equivalent to $\operatorname{Tr}(T)=2\cos(K)$ and the above mentioned condition simplifies to
$|\mathrm{Tr}(T)|\leq 2\quad\implies\quad\text{bulk mode}.$ (7)
For an edge mode, since the boundary condition imposes $a_{0}=0$, the $T$-matrix must instead satisfy
$T\begin{pmatrix}0\\ 1\end{pmatrix}=\alpha\begin{pmatrix}0\\ 1\end{pmatrix}{,\ \ \ }\alpha\in(-1,1)\implies\text{edge mode}.$ (8)
The condition $|\alpha|<1$ ensures the mode is localized at the edge. Notably, when Eq. (8) is satisfied, we automatically have $|\mathrm{Tr}(T)|>2,$ so a $T$ matrix cannot support both an edge mode and a bulk mode at the same energy $E$.
@@ III Topological Protection in the Transfer Matrix Formalism
Figure 1: (a) Illustration of the SSH model. The open circles denote sites at which the electron can sit, and the lines denote the coupling strength between each site. In the SSH model, the couplings alternate between two values $\gamma_{1}$ and $\gamma_{2}$. (b) The band structure of the SSH model $(\mu_{1}=\mu_{2})$ embedded in $\mathrm{SL}(2,\mathbb{R})$ (3D Iwasawa parameters space). The parameter $k$ is only defined modulo $2\pi$. The plot is then unfolded in the same way we conventionally unfold band diagrams beyond the first Brillouin zone. Each curve represents a Hamiltonian, or equivalently, a family of $T$-matrices $T(E)$ for a different value of $\gamma_{1}/\gamma_{2}$. When the curve touches one of the orange surfaces, an edge state is supported. The decay length of this edge state is given by $|\ln(a)|^{-1}$. The green regions are the allowed bands: Each point on the curve inside the green region corresponds to a bulk propagating mode. The thick black lines on the boundary of the orange surface are the points we want to exclude so that our model to be topologically nontrivial.
@@ III.1 The $\mathbb{R}$) Picture for Topological Phases
$T=\begin{pmatrix}\cos(k)&\sin(k)\\ -\sin(k)&\cos(k)\end{pmatrix}\begin{pmatrix}a^{-1}&0\\ 0&a\end{pmatrix}\begin{pmatrix}1&0\\ n&1\end{pmatrix},$ (9)
where $k,n\in\mathbb{R},a>0$ are the Iwasawa parameters, and they are smooth functions of the entries $T_{ij}$. By doing so, we can parameterise $T(E)$ as a curve in a three-dimensional space, where the coordinates are taken to be the Iwasawa parameters: $k(E)$, $a(E)$, $n(E)$. We plot a family of curves with various choices of $\gamma$ in Fig. 1 for the SSH model. Since $\det T\in\mathrm{SL}(2,\mathbb{R})$ the condition that the bulk states must stay bounded in the chain is automatically satisfied and we may identify the transfer matrices supporting bulk states by simply looking at the condition $\mathrm{Tr}(T)\leq 2$. On the other hand for a $T$ matrix to support an edge state we follow condition imposed by Eq. (8): we require $(0,1)$ to be an eigenvector of $T$ with eigenvalue $\alpha$ less than one, $|\alpha|<1$. In terms of the Iwasawa parameters, this imposes $n=0$, $k\in\pi\mathbb{Z}$, and $a<1$. This way, $\alpha=\cos(k)a=\pm a$.
@@
Topological insulators are characterised by edge states which cannot disappear unless the material passes through a conducting state [27]. In the above picture, this means that we need to somehow exclude all the points $k=\pi\mathbb{Z},\ln(a)=0,n\neq 0$ (these are the thick black lines in 1). Any kind of constraint on the Hamiltonian or $T$-matrix which excludes these points will do the job just fine, but this is usually done using symmetries. In agreement with Zak’s argument [28], parity symmetry automatically gives us the following relation for the Iwasawa parameters
$\sigma_{x}T\sigma_{x}=T^{-1}\quad\implies\quad n=\left(\frac{1}{a^{2}}-1\right)\tan(k),$ (10)
@@ III.2 Proof of the Zak Phase Bulk-Edge Correspondence
Here, we show how our formalism so far is connected to Bloch band topological invariants. In doing so, we will obtain a rigorous but pictorially simple proof of the bulk-edge correspondence. We stress that our results hold for any number of sites per unit cell, as well as for continuous quantum mechanical or photonic systems. This will become more clear in Sec. IV.1 where we treat the case of $4$ sites per unit cell more explicitly. But for now, we will simply highlight a few key features when we increase the number of sites per unit cell.
@@
where $\mathcal{N}$ is a normalisation constant, $\phi(K)$ is a real function, and the entries in the dots are not important (see Appendix A for details).
@@
$\phi_{\text{Zak}}\equiv\frac{1}{\mathrm{i}\mkern 1.0mu}\int\limits_{0}^{2\pi}\left\langle K\right|\partial_{K}\left|K\right\rangle\mathop{}\!\mathrm{d}K=\frac{\phi(2\pi)-\phi(0)}{2},$ (12)
where $\left|K\right\rangle$ is the Bloch eigenvector of the system. This result is well-known for the case of 2 sites per unit cell [20]. We provide details of the general case in Appendix A.
@@
$\phi(K)=K-\varphi(K),$ (13)
where $\varphi(K)$ is the relative phase between the two entries of the $T$-eigenvector with eigenvalue $\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0muK}$ (Fig. 2). At the same time, this can be interpreted as the relative phase between the last element of the Bloch eigenvector $\left|K\right\rangle$ and the first element $\left|K\right\rangle\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0muK}$ of the next unit cell.
Figure 2: Illustration of the connection between the transfer matrix picture and the Bloch Hamiltonian picture. The $T$-matrix connects only the leftmost and rightmost nodes of unit cells.
It should be noted here that both $\phi$ and $\varphi$ depend on the choice of basis. This is why the middle expression in (13) must be written in a basis-dependent way. This basis dependence will not be a problem though since in the end we are only interested in differences $\phi(2\pi)-\phi(0)$ and $\varphi(2\pi)-\varphi(0)$, where the basis dependence will cancel.
The parity symmetry condition (10) along with the condition $|\mathrm{Tr}(T)|=2$, that the Bloch eigenvector is $K\equiv 0\pmod{\pi}$, fixes $T$ to lie in one of four branches of solutions (See Appendix B). As an example, one of the possible solutions is given by
$T=\left(\begin{array}[]{cc}-\frac{\cos(k)}{\sin(k)+\cos(k)}&\frac{\sin(k)}{\sin(k)+\cos(k)}\\ -\frac{\sin(k)}{\sin(k)+\cos(k)}&-\frac{2\sin(k)+\cos(k)}{\sin(k)+\cos(k)}\\ \end{array}\right),$ (14)
@@
Figure 3: Bulk-edge correspondence for the Zak phase, as viewed in $\mathbb{R}$), and projected to $n$ given by Eq. (10). For clarity we have included example $T(E)$ paths using the SSH4 model described in (24). Note that the parameter $k$ is only defined modulo $2\pi$. The plot has been unfolded in much the same way as a band diagram can be unfolded beyond the first Brillouin zone. The green regions denote transfer matrices which support bulk modes (allowed bands regions), while the brown lines denote transfer matrices which support edge modes. The purple dashed boundaries denote the points where $\phi=\pi$, and the red dashed curves denote the points where $\phi=0$. The Zak phase $\phi_{\text{Zak}}$ is $\pi$ when the $T(E)$ curve enters and exits from boundaries of different colours, and it is 0 when the two colours are the same. As such, we see that the creation of edge states (magenta stars) is necessarily accompanied by the change in Zak phase of the two neighbouring bands.
This shows that in general the number of edge state in the forbidden zone above the $j^{\text{th}}$ band is the sum of Zak phases of all the lower bands, modulo 2: